OpenStudy (anonymous):

How do I set these two equations equal to each other? 16=x+y 8=x-y

OpenStudy (anonymous):

x = 8 y = 8

OpenStudy (anonymous):

That doesn't make sense.

OpenStudy (midhun.madhu1987):

multiply the second equation by 2

OpenStudy (midhun.madhu1987):

what u get?

OpenStudy (bibby):

8=x-y Multiply both sides by 2 16 = 2(x-y)

OpenStudy (anonymous):

Oops. I did it wrong. 1 sec.

OpenStudy (anonymous):

x = 12 y = 4

OpenStudy (anonymous):

Should x be 12 and y is 4. Yeah. Okay.

OpenStudy (anonymous):

Thanks. :)

OpenStudy (anonymous):

No problem.

OpenStudy (bibby):

http://openstudy.com/code-of-conduct